OpenStudy (anonymous):

***(y-k)=a(x-h)^2 . Find the x-intercept of parabola with vertex(-5,80) and y intercept (0,-45) in (x1,y1) ( x2,y2) form :D

myininaya (myininaya):

y=a(x-h)^2+k y=a(x+5)^2+80 -45=a(0+5)^2+80 -45-80=a*25 -125=25a a=-125/25=-5 so we have y=-5(x+5)^2+80

OpenStudy (anonymous):

see? told you someone would help :)